# Cross-Layer Vision Smoothing: Enhancing Visual Understanding via Sustained Focus on Key Objects in Large Vision-Language Models

### v1

@@ Abstract
Large Vision-Language Models (LVLMs) can accurately locate key objects in images, yet their attention to these objects tends to be very brief. Motivated by the hypothesis that sustained focus on key objects can improve LVLMs’ visual capabilities, we propose Cross-Layer Vision Smoothing (CLVS). The core idea of CLVS is to incorporate a vision memory that smooths the attention distribution across layers. Specifically, we initialize this vision memory with position-unbiased visual attention in the first layer. In subsequent layers, the model’s visual attention jointly considers the vision memory from previous layers, while the memory is updated iteratively, thereby maintaining smooth attention on key objects. Given that visual understanding primarily occurs in the early and middle layers of the model, we use uncertainty as an indicator of completed visual understanding and terminate the smoothing process accordingly. Experiments on four benchmarks across three LVLMs confirm the effectiveness and generalizability of our method. CLVS achieves state-of-the-art performance on a variety of visual understanding tasks, with particularly significant improvements in relation and attribute understanding. [^1] Our code is available at https://github.com/beta-nlp/CLVS.
@@ 1 Introduction
Large Vision-Language Models (LVLMs) (Liu et al. 2023; Bai et al. 2025) acquire the capability to comprehend visual information through training on large-scale multimodal data and exhibit strong performance across a wide range of visual understanding tasks (Li et al. 2023; Wang et al. 2023; Fu et al. 2024). Despite these successes, LVLMs still suffer from issues such as hallucinations (Leng et al. 2024), which limit their real-world applications.
Recent studies (Zhang et al. 2025; Chen et al. 2025) have revealed a strong correlation between the visual understanding capability of LVLMs and their visual attention. Visual attention is learned through end-to-end training without direct supervision. While this approach simplifies LVLM training, it also makes it difficult to avoid biases in visual attention (Liu, Zheng, and Chen 2024; Kang et al. 2025; Zhu et al. 2025b), such as position bias (Zhu et al. 2025b; Li et al. 2025), visual attention sink (Huang et al. 2024; Kang et al. 2025), and uncoordinated attention heads (He et al. 2025; Yang et al. 2025). Consequently, a series of recent works have sought to enhance the visual understanding capabilities of LVLMs by improving visual attention. Liu, Zheng, and Chen (2024); Zhu et al. (2025a); Yin, Si, and Wang (2025a) propose directly increasing the attention weights on visual tokens. Yang et al. (2025); He et al. (2025) analyze the effect of the multi-head attention mechanism on visual understanding and remove attention heads that negatively impact performance. Kang et al. (2025); Xie et al. (2025); Chen et al. (2025) adjust the visual attention distribution to achieve finer control over the model’s visual understanding process. These successes confirm the crucial role of visual attention in the visual understanding of LVLMs.
Figure 1: A running example of visual attention in LLaVA-1.5-7B. In vanilla LVLMs, first-layer attention is biased toward the bottom, and focus on the key object (woman) decays after layer 15.
@@
While previous studies have demonstrated effectiveness in improving the visual attention of LVLMs, they primarily focus on enhancing attention within individual layers without considering the cross-layer evolution of visual attention. We observe that visual attention can change dramatically between consecutive layers. For example, in Figure 1, when asked “Is the woman wearing a black shirt?”, the model attends to the “woman” and “shirt” (i.e., key objects) at layer 14 but ceases to focus on them afterward. Although the model accurately locates key objects, its attention to these objects is very brief. This phenomenon, commonly observed across different LVLMs and images, is referred to as advantageous attention decay. We hypothesize that such brief attention is insufficient for effective visual information fusion in LVLMs. Therefore, sustaining an appropriate focus on key objects is beneficial to enhance the model’s visual capabilities.
To tackle advantageous attention decay, we propose Cross-Layer Vision Smoothing (CLVS) to enhance LVLMs’ visual comprehension by maintaining sustained focus on key objects in the image. To this end, we introduce a vision memory for smoothing visual attention across layers. Specifically, following Li et al. (2025), we unify the visual attention in the first layer for unbiased perception and initialize the vision memory with the visual attention distribution. In subsequent layers, we progressively smooth visual attention by interpolating it with the vision memory and updating the memory iteratively. Moreover, since visual understanding primarily occurs in the early and middle layers of LVLMs, we propose an uncertainty-based criterion to determine whether the visual understanding process has been completed and, if so, terminate visual attention smoothing. With CLVS, the model can sustain focus on key objects, thereby improving the visual capabilities of LVLMs. The comparison of visual attention to key objects across layers between LVLMs with and without CLVS is shown in Figure 2.
We conduct extensive experiments on four visual understanding benchmarks and three LVLMs. The results demonstrate that CLVS effectively improves the visual capabilities of LVLMs, reducing hallucinations, particularly those related to object attributes and relations.
@@ 2 Related Work
The cross-modal understanding ability of LVLMs is built upon the attention mechanism. However, end-to-end training lacks supervision over the model’s internal processes, leading to shortcomings in the spontaneously formed visual attention, such as position bias and visual attention sink. These issues can result in hallucinations (Liu et al. 2024; Zhao et al. 2025b) in the model’s output.
Numerous works have explored improving visual attention in LVLMs. Yang et al. (2025); Zhu et al. (2025b) incorporate supervisory signals into the attention mechanism, enabling fine-grained control during training. To avoid high training costs and improve generalizability, an increasing number of studies focus on training-free approaches to optimize visual attention. Given the critical role of visual information in LVLMs, Liu, Zheng, and Chen (2024); Zhu et al. (2025a); Yin, Si, and Wang (2025a) directly increase the model’s attention weights on visual tokens. While effective, these approaches overlook the internal dynamics of visual attention. Leveraging the diversity of multi-head attention, Zhang et al. (2024); Zhao et al. (2025a); Yang et al. (2025) identify and utilize high-quality attention heads to optimize visual attention. Recent studies (Kang et al. 2025; Xie et al. 2025) further analyze attention distributions to identify the causes of hallucinations and implement targeted optimizations. However, these works enhance visual attention independently within each layer, overlooking its evolution across layers.
@@ 3 Cross-Layer Vision Smoothing
Figure 3: Overview of CLVS. S, V, U, and G denote system tokens, visual tokens, user tokens, and the current generating token, respectively. Unified position indices are applied in the first layer. The vision memory, initialized from this unbiased visual attention, smooths and is updated across subsequent layers.
@@ 3.1 Unified Visual Positions
Positional embeddings are determined by the absolute or relative position of a token in a sequence and play an important role in Transformer-based models. Recent studies (Zhu et al. 2025b; Li et al. 2025) have found that Transformer-based LVLMs tend to focus disproportionately on the bottom-right corner of the image (i.e., the visual tokens at the end of the input sequence), a bias influenced by the positions of visual tokens in the inputs. Unlike textual inputs, visual information does not inherently possess sequential characteristics. Therefore, the model should receive unbiased attention, particularly in the first layer, where it has no prior visual perception.
Following Li et al. (2025), we normalize the positional indices of all image tokens to a unified index and replace the original positional indices in the concatenated input sequence. Given a sequence of length $N=N_{s}+N_{v}+N_{i}$, the new positional indices are formally defined as
@@
where $N_{s}$, $N_{v}$, and $N_{i}$ denote the number of tokens in the system prompt, image, and user prompt, respectively. Note that since we only adjust the visual attention of the current generating token, the unified position does not affect the attention among visual tokens. Because of the causal mask in the attention layer, information flows unidirectionally through the input sequence (Zhao et al. 2025a; Yin, Si, and Wang 2025b), which aligns with the positional indices. Therefore, we apply position debiasing to the input image only in the first layer.
@@ 3.2 Visual Attention Smoothing
Previous studies (Zhang et al. 2025; Chen et al. 2025) reveal a strong correlation between the visual understanding capabilities of LVLMs and their visual attention. However, we found that the model’s attention quickly converges on sink tokens, hindering sustained focus on key objects. To mitigate this issue and ensure stable attention on key objects, we introduce a vision memory $\bm{m}$ into the LVLM, which serves as a cross-layer smoothing mechanism in visual attention. Let $\bm{\alpha}^{(l)}_{h}=\{\bm{\lambda}^{(l)}_{h},\bm{\mu}^{(l)}_{h}\}$, where $\bm{\lambda}^{(l)}_{h}\in\mathbb{R}^{N_{v}}$ denotes the attention weights associated with visual tokens ($N_{v}$ is the number of visual tokens), and $\bm{\mu}^{(l)}_{h}\in\mathbb{R}^{N-N_{v}}$ denotes the attention weights associated with non-visual tokens. The vision memory $\bm{m}$ stores only the attention weights corresponding to visual tokens. We initialize the vision memory at the first layer, denoted as $\bm{m}^{(1)}$, using the unbiased visual attention of that layer, and apply max pooling across attention heads to compress the model’s overall attention to the image:
@@ 3.3 Vision Understanding Completion
The visual understanding process in LVLMs primarily occurs in the early and middle layers, while language-level feature construction takes place in the final layers (Yin, Si, and Wang 2025b; Zhao et al. 2025a). Therefore, adjustments to visual attention should target the early and middle layers. Previous works (Xie et al. 2025; Yin, Si, and Wang 2025a) often determine the layer range for attention optimization by setting fixed hyperparameters, which limits generalization across different LVLMs. To address this, we propose a dynamic mechanism that determines the active layer range by assessing whether the model’s visual understanding process is complete, and terminates CLVS accordingly.
Intuitively, if the model’s output has converged (i.e., the output token remains unchanged in subsequent layers under greedy decoding), we can terminate the smoothing process. However, output convergence is determined based on the final layer’s output and therefore cannot be known at the current layer. A straightforward approach would be to perform a preliminary forward pass to determine the convergence status at each layer, but this significantly increases inference cost. Following Zou et al. (2025), we instead use output uncertainty at each layer as a surrogate measure to determine the end of visual understanding. Specifically, we multiply the representation by the unembedding matrix to obtain the probability distribution at the current layer $l$, and calculate the uncertainty measure $u$ as
@@
where $|V|$ denotes the vocabulary size. For computational concerns, we follow the practice of Zou et al. (2025) and compute only the top-$10$ tokens (i.e., $|V|=10$).
Figure 4: Statistical results of output uncertainty and convergence on the MME benchmark. Left: clear inverse relationship between convergence and uncertainty. Right: “PB Correlation” denotes Point-Biserial Correlation, and “LR Coefficient” denotes the coefficient from Logistic Regression analysis.
We analyze the uncertainty and convergence of the model’s responses, with results shown in Figure 4. From the left panel, we observe a strong correlation between convergence and uncertainty: as uncertainty decreases, the model progressively completes visual perception and begins constructing semantic information. Both point-biserial correlation and logistic regression analyses, shown in the right panel, further confirm statistical associations between these measures. Based on these observations, we use uncertainty to determine whether the model has completed visual perception. Since this process is completed by at least the middle layers, we introduce uncertainty detection starting from the $(L+1)/2$-th layer to avoid unnecessary computations. Visual attention smoothing is terminated in subsequent Transformer layers once the uncertainty falls below the threshold $\delta$.
Model Method Exsitence Attribute Relation All
Acc. F1 Acc. F1 Acc. F1 Acc. F1
LLaVA-1.5-7B Vanilla 71.4 83.3 72.0 64.4 74.1 68.5 72.1 74.8
VCD (Leng et al. 2024) 68.3 81.1 72.7 66.6 72.4 65.1 71.1 74.1
PAI (Liu, Zheng, and Chen 2024) 75.7 86.1 71.9 64.2 75.8 71.2 73.7 76.7
VAR (Kang et al. 2025) 70.7 82.8 72.1 65.0 73.7 66.6 71.8 74.6
AD-HH (Yang et al. 2025) 71.4 83.3 71.8 64.2 74.0 68.7 71.9 74.8
VAF  (Yin, Si, and Wang 2025a) 64.4 78.3 70.4 61.7 73.3 65.3 68.7 70.9
TARAC (Xie et al. 2025) 72.1 83.7 71.4 63.5 74.2 68.4 72.0 74.8
CLVS 77.6 87.3 72.7 65.4 77.1 73.7 75.0 78.1
LLaVA-1.5-13B Vanilla 66.0 79.5 75.3 69.5 70.1 45.2 71.5 73.5
VCD (Leng et al. 2024) 64.6 78.4 75.9 71.3 69.9 45.7 71.3 73.5
PAI  (Liu, Zheng, and Chen 2024) 70.5 82.6 78.1 74.3 73.9 56.6 75.0 77.6
VAR  (Kang et al. 2025) 65.2 78.9 75.1 69.1 69.7 43.8 71.0 72.9
AD-HH (Yang et al. 2025) 67.1 80.3 75.8 70.4 70.0 45.5 72.1 74.2
VAF  (Yin, Si, and Wang 2025a) 61.4 76.0 74.9 68.9 69.4 43.5 69.5 71.2
TARAC (Xie et al. 2025) 68.7 81.4 76.2 70.9 71.3 48.7 73.0 75.3
CLVS 78.4 87.8 78.9 76.7 76.2 64.3 78.4 81.8
@@
VCD (Leng et al. 2024) 95.0 97.4 84.4 85.2 67.7 71.8 86.1 89.9
VAF (Yin, Si, and Wang 2025a) 93.0 96.3 84.7 85.3 71.5 74.2 86.0 89.7
TARAC (Xie et al. 2025) 95.0 97.4 84.2 85.1 66.2 70.9 85.8 89.8
CLVS 93.5 96.6 85.2 85.7 72.1 74.5 86.6 90.0
Qwen2.5-VL Vanilla 95.3 97.5 83.5 84.4 73.9 75.8 86.5 90.1
VCD  (Leng et al. 2024) 93.0 96.3 83.7 84.3 73.3 75.0 85.7 89.4
VAF (Yin, Si, and Wang 2025a) 93.8 96.8 83.7 84.5 71.5 74.1 85.7 89.5
TARAC (Xie et al. 2025) 95.4 97.6 83.5 84.5 73.6 75.6 86.5 90.1
CLVS 94.8 97.3 83.9 84.8 74.6 76.2 86.6 90.2

### v2

@@ Abstract
Large Vision-Language Models (LVLMs) can accurately locate key objects in images, yet their attention to these objects tends to be very brief. Motivated by the hypothesis that sustained focus on key objects can improve LVLMs’ visual capabilities, we propose Cross-Layer Vision Smoothing (CLVS). The core idea of CLVS is to incorporate a vision memory that smooths the attention distribution across layers. Specifically, we initialize this vision memory with position-unbiased visual attention in the first layer. In subsequent layers, the model’s visual attention jointly considers the vision memory from previous layers, while the memory is updated iteratively, thereby maintaining smooth attention on key objects. Given that visual understanding primarily occurs in the early and middle layers of the model, we use uncertainty as an indicator of completed visual understanding and terminate the smoothing process accordingly. Experiments on four benchmarks across three LVLMs confirm the effectiveness and generalizability of our method. CLVS achieves state-of-the-art overall performance across a variety of visual understanding tasks and attains comparable results to the leading approaches on image captioning benchmarks. [^1] The code is available at: https://github.com/beta-nlp/CLVS.
@@ 1 Introduction
Large Vision-Language Models (LVLMs) [llava, qwen25-vl] acquire the capability to comprehend visual information through training on large-scale multimodal data and exhibit strong performance across a wide range of visual understanding tasks [POPE, AMBER, MME]. Despite these successes, LVLMs still suffer from issues such as hallucinations [VCD], which limit their real-world applications.
Recent studies [ViCrop, AdaptVis] have revealed a strong correlation between the visual understanding capability of LVLMs and their visual attention. Visual attention is learned through end-to-end training without direct supervision. While this approach simplifies LVLM training, it also makes it difficult to avoid biases in visual attention [PAI, VAR, DAC], such as position bias [DAC, IMCCD], visual attention sink [opera, VAR], and uncoordinated attention heads [VHD, AD-HH]. Consequently, a series of recent works have sought to enhance the visual understanding capabilities of LVLMs by improving visual attention. Some works [PAI, IBD, VAF] propose directly increasing the attention weights on visual tokens. other works [AD-HH, VHD] analyze the effect of the multi-head attention mechanism on visual understanding and remove attention heads that negatively impact performance. Further, several methods [VAR, TARAC, AdaptVis] adjust the visual attention distribution to achieve finer control over the model’s visual understanding process. These successes confirm the crucial role of visual attention in the visual understanding of LVLMs.
Figure 1: A running example of visual attention in LLaVA-7B. In vanilla LVLMs, first-layer attention is biased toward the bottom, and focus on the key object (woman) decays after layer 14.
@@
While previous studies have demonstrated effectiveness in improving the visual attention of LVLMs, they primarily focus on enhancing attention within individual layers without considering the cross-layer evolution of visual attention. We observe that visual attention can change dramatically between consecutive layers. For example, in Fig. 1, when asked “Is the woman wearing a black shirt?”, the model attends to the “woman” and “shir” (i.e., key objects) at layer 14 but ceases to focus on them afterward. Although the model accurately locates key objects, its attention to these objects is very brief. This phenomenon, commonly observed across different LVLMs and images, is referred to as advantageous attention decay. We hypothesize that such brief attention is insufficient for effective visual information fusion in LVLMs. Therefore, sustaining an appropriate focus on key objects is beneficial to enhance the model’s visual capabilities.
To tackle advantageous attention decay, we propose Cross-Layer Vision Smoothing (CLVS) to enhance LVLMs’ visual comprehension by maintaining sustained focus on key objects in the image. To this end, we introduce a vision memory for smoothing visual attention across layers. Specifically, following previous work [IMCCD], we unify the visual attention in the first layer for unbiased perception and initialize the vision memory with the visual attention distribution. In subsequent layers, we progressively smooth visual attention by interpolating it with the vision memory and updating the memory iteratively. Moreover, since visual understanding primarily occurs in the early and middle layers of LVLMs, we propose an uncertainty-based criterion to determine whether the visual understanding process has been completed and, if so, terminate visual attention smoothing. With CLVS, the model can sustain focus on key objects, thereby improving the visual capabilities of LVLMs. The comparison of visual attention to key objects across layers between LVLMs with and without CLVS is shown in Fig. 2.
@@ 2 Related Work
The cross-modal understanding ability of LVLMs is built upon the attention mechanism. However, end-to-end training lacks supervision over the model’s internal processes, leading to shortcomings in the spontaneously formed visual attention, such as position bias and visual attention sink. These issues can result in hallucinations [hallu_lvlm, CICD] in the model’s output.
Numerous works have explored improving visual attention in LVLMs. AD-HH, DAC incorporate supervisory signals into the attention mechanism, enabling fine-grained control during training. To avoid high training costs and improve generalizability, an increasing number of studies focus on training-free approaches to optimize visual attention. Given the critical role of visual information in LVLMs, PAI, IBD, VAF directly increase the model’s attention weights on visual tokens. While effective, these approaches overlook the internal dynamics of visual attention. Leveraging the diversity of multi-head attention, EAH, SEVI, AD-HH identify and utilize high-quality attention heads to optimize visual attention. Recent studies [VAR, TARAC] further analyze attention distributions to identify the causes of hallucinations and implement targeted optimizations. However, these works enhance visual attention independently within each layer, overlooking its evolution across layers.
@@ 3 Cross-Layer Vision Smoothing
Figure 3: Overview of CLVS. S, V, and T denote system tokens, visual tokens, and textual tokens respectively. Unified position indices are applied in the first layer. The vision memory, initialized from this unbiased visual attention, smooths and is updated across subsequent layers.
@@ 3.1 Unified Visual Positions
Positional embeddings are determined by the absolute or relative position of a token in a sequence and play an important role in Transformer-based models. Recent studies [DAC, IMCCD] have found that Transformer-based LVLMs tend to focus disproportionately on the bottom-right corner of the image (i.e., the visual tokens at the end of the input sequence), a bias influenced by the positions of visual tokens in the inputs. Unlike textual inputs, visual information does not inherently possess sequential characteristics. Therefore, the model should receive unbiased attention, particularly in the first layer, where it has no prior visual perception.
Following previous work [IMCCD], we normalize the positional indices of all image tokens to a unified index and replace the original positional indices in the concatenated input sequence. Given a sequence of length $N=N_{s}+N_{v}+N_{i}$, the new positional indices are formally defined as
@@
where $N_{s}$, $N_{v}$, and $N_{i}$ denote the number of tokens in the system prompt, image, and user prompt, respectively. Note that since we only adjust the visual attention of the current generating token, the unified position does not affect the attention among visual tokens. Because of the causal mask in the attention layer, information flows unidirectionally through the input sequence [SEVI, VisInfoFlow], which aligns with the positional indices. Therefore, we apply position debiasing to the input image only in the first layer.
@@ 3.2 Visual Attention Smoothing
Previous studies [ViCrop, AdaptVis] reveal a strong correlation between the visual understanding capabilities of LVLMs and their visual attention. However, we found that the model’s attention quickly converges on sink tokens, hindering sustained focus on key objects. To mitigate this issue and ensure stable attention on key objects, we introduce a vision memory $\bm{m}$ into the LVLM, which serves as a cross-layer smoothing mechanism in visual attention. Let $\bm{\alpha}^{(l)}_{h}=\{\bm{\lambda}^{(l)}_{h},\bm{\mu}^{(l)}_{h}\}$, where $\bm{\lambda}^{(l)}_{h}\in\mathbb{R}^{N_{v}}$ denotes the attention weights associated with visual tokens ($N_{v}$ is the number of visual tokens), and $\bm{\mu}^{(l)}_{h}\in\mathbb{R}^{N-N_{v}}$ denotes the attention weights associated with non-visual tokens. The vision memory $\bm{m}$ stores only the attention weights corresponding to visual tokens. We initialize the vision memory at the first layer, denoted as $\bm{m}^{(1)}$, using the unbiased visual attention of that layer, and apply max pooling across attention heads to compress the model’s overall attention to the image:
@@ 3.3 Vision Understanding Completion
The visual understanding process in LVLMs primarily occurs in the early and middle layers, while language-level feature construction takes place in the final layers [VisInfoFlow, SEVI]. Therefore, adjustments to visual attention should target the early and middle layers. Previous works [TARAC, VAF] often determine the layer range for attention optimization by setting fixed hyperparameters, which limits generalization across different LVLMs. To address this, we propose a dynamic mechanism that determines the active layer range by assessing whether the model’s visual understanding process is complete, and terminates CLVS accordingly.
Intuitively, if the model’s output has converged (i.e., the output token remains unchanged in subsequent layers under greedy decoding), we can terminate the smoothing process. However, output convergence is determined based on the final layer’s output and therefore cannot be known at the current layer. A straightforward approach would be to perform a preliminary forward pass to determine the convergence status at each layer, but this significantly increases inference cost. We instead use output uncertainty [LookTwice] at each layer as a surrogate measure to determine the end of visual understanding. Specifically, we multiply the representation by the unembedding matrix to obtain the probability distribution at the current layer $l$, and calculate the uncertainty measure $u$ as
@@
where $|V|$ denotes the vocabulary size. For computational concerns, we compute only the top-$10$ tokens (i.e., $|V|=10$).
Figure 4: Statistical results of output uncertainty and convergence on the MME benchmark. Left: clear inverse relationship between convergence and uncertainty. Right: PB Correlation denotes Point-Biserial Correlation, and LR Coefficient denotes the coefficient from Logistic Regression analysis.
We analyze the uncertainty and convergence of the model’s responses, with results shown in Fig. 4. From the left panel, we observe a strong correlation between convergence and uncertainty: as uncertainty decreases, the model progressively completes visual perception and begins constructing semantic information. Both point-biserial correlation and logistic regression analyses, shown in the right panel, further confirm statistical associations between these measures. Based on these observations, we use uncertainty to determine whether the model has completed visual perception. Once the uncertainty falls below the threshold $\delta$, Visual attention smoothing is terminated in subsequent Transformer layers.
Model Method Existence Attribute Relation All
Acc. $\uparrow$ F1 $\uparrow$ Acc. $\uparrow$ F1 $\uparrow$ Acc. $\uparrow$ F1 $\uparrow$ Acc. $\uparrow$ F1 $\uparrow$
LLaVA-7B Vanilla 71.4 83.3 72.0 64.4 74.1 68.5 72.1 74.8
VCD 68.3 81.1 72.7 66.6 72.4 65.1 71.1 74.1
PAI 70.3 82.5 69.5 59.1 74.5 65.6 70.3 72.5
PAICD 75.7 86.1 71.9 64.2 75.8 71.2 73.7 76.7
VAR 70.7 82.8 72.1 65.0 73.7 66.6 71.8 74.6
AD-$H$ 71.4 83.3 71.8 64.2 74.0 68.7 71.9 74.8
TARAC 72.1 83.7 71.4 63.5 74.2 68.4 72.0 74.8
CLVS 80.1 88.9 73.7 67.3 76.7 74.9 76.2 79.6
LLaVA-13B Vanilla 66.0 79.5 75.3 69.5 70.1 45.2 71.5 73.5
VCD 64.6 78.4 75.9 71.3 69.9 45.7 71.3 73.5
PAI 69.2 81.7 77.5 73.2 73.9 56.7 74.2 76.7
PAICD 70.5 82.6 78.1 74.3 73.9 56.6 75.0 77.6
VAR 65.2 78.9 75.1 69.1 69.7 43.8 71.0 72.9
AD-$H$ 67.1 80.3 75.8 70.4 70.0 45.5 72.1 74.2
TARAC 68.7 81.4 76.2 70.9 71.3 48.7 73.0 75.3
CLVS 76.3 86.5 78.7 75.1 74.9 59.9 77.4 80.3
@@
VCD [VCD] 95.0 97.4 84.4 85.2 67.7 71.8 86.1 89.9
PAI 94.9 97.3 84.6 85.3 70.0 73.2 86.4 90.1
TARAC [TARAC] 95.0 97.4 84.2 85.1 66.2 70.9 85.8 89.8
CLVS 94.3 97.0 85.1 85.6 70.3 73.3 86.6 90.1
Table 2: Results on POPE and R-Bench. POPE’s results are the average performance across the MSCOCO, A-OKVQA, and GQA datasets.
Model Method Random Popular Adversarial R-Bench
[…]
LLaVA-7B Vanilla 87.12 87.94 79.38 82.14 71.85 77.04 67.76 74.84
VCD 84.89 86.07 77.38 80.64 71.03 76.39 68.22 74.97
PAI 86.20 87.20 78.45 81.56 70.83 76.46 67.17 74.66
PAICD 87.27 87.98 80.03 82.51 72.61 77.38 68.69 75.39
AD-$H$ 87.13 87.93 79.43 82.18 71.89 77.07 67.60 74.76
VAF 84.88 86.34 77.13 80.82 69.59 75.94 66.79 74.45
TARAC 86.99 87.80 79.37 82.10 71.92 77.03 67.67 74.83
CLVS 88.74 88.98 82.17 83.76 75.34 78.80 69.01 75.47
LLaVA-13B Vanilla 84.31 85.96 78.64 81.87 71.58 77.23 68.18 75.39
VCD 83.01 84.81 77.37 80.81 71.38 76.84 68.88 75.48
PAI 85.36 86.71 79.72 82.56 72.38 77.64 70.31 76.51
PAICD 86.10 87.24 80.86 83.31 73.44 78.26 71.00 76.89
AD-$H$ 84.88 86.37 79.28 82.29 72.26 77.62 68.61 75.61
VAF 82.42 84.62 77.10 80.91 70.08 76.41 67.66 75.10
TARAC 87.02 87.80 79.40 82.10 71.95 77.03 68.79 75.71
CLVS 87.26 88.08 82.73 84.55 75.09 79.16 70.80 76.79
LLaVA-Next Vanilla 87.36 85.90 85.45 84.14 82.31 81.40 81.20 81.88
PAI 87.71 86.34 85.95 84.70 82.64 81.82 81.88 82.57
TARAC 86.13 84.26 83.94 82.25 81.04 79.73 81.76 82.33
CLVS 88.69 87.59 86.21 85.31 82.80 82.36 81.67 82.60
Table 3: Results on image captioning task.
Model Method AMBER CHAIR
CHAIR $\downarrow$ Cover $\uparrow$ Hal $\downarrow$ Cog $\downarrow$ CHAIRs $\downarrow$ CHAIRi $\downarrow$ F1 $\uparrow$
LLaVA-7B Vanilla 7.1 50.7 32.5 3.8 52.2 13.9 76.1
PAI 5.8 48.8 27.5 2.5 33.6 9.0 77.1
PAICD 5.1 48.4 25.7 2.1 30.8 8.1 76.3
TARAC 5.8 49.3 28.5 2.9 43.0 11.2 75.1
CLVS 4.2 48.0 21.9 2.1 33.0 9.2 75.9
LLaVA-13B Vanilla 6.8 51.3 31.1 3.4 52.6 14.1 76.4
PAI 5.4 51.6 27.3 2.1 40.2 10.8 76.9
PAICD 5.2 52.0 28.3 1.9 37.0 9.9 77.3
TARAC 5.8 49.3 28.5 2.9 43.0 11.2 75.1
CLVS 4.9 50.9 23.7 2.0 43.8 11.5 77.0
LLaVA-Next Vanilla 7.8 63.1 46.1 3.9 34.2 9.3 72.2
PAI 7.1 64.2 42.2 3.2 34.0 8.9 73.7
TARAC 6.7 60.3 37.5 3.4 29.8 8.0 71.5
CLVS 7.1 62.4 42.9 3.4 32.2 8.2 71.7